OpenStudy (anonymous):

Vinegar has a density of 1.0056 g/cm3. What is the mass of 7 L of vinegar?

OpenStudy (frostbite):

Solve the problem in two steps: 1) rewrite the density to the unit gram/L 2 ) \[\large \rho = \frac{ m }{ V } \to m=\rho ~ V\]

OpenStudy (frostbite):

rho (the p like thing) is the density. m the mass V the volume

OpenStudy (anonymous):

I did all that and got 7039.2 am I doing something wrong

OpenStudy (frostbite):

Looks right to me: \[\large 1.0056 \frac{ g }{ cm ^{3} }*1000 \frac{ cm ^{3} }{ dm ^{3} }=1005.6 \frac{ g }{ dm ^{3} }=1005.6 \frac{ g }{ L }\] \[\large 1005.6 \frac{ g }{ L } * 7 L \approx 7039 g\]

OpenStudy (anonymous):

I put that into my webassign homework and it marks it wrong, I tried 7039.2, 7039, and 7040

OpenStudy (frostbite):

Okay. Funny, but I'm 100% sure on my calculation, the conversion should also be right: Lets go it through: \[\large 1 dm ^{3}=1 L\] This is per definition. \[\large 1000 ~ cm ^{3}=1 ~ dm ^{3} \to 1=1000 ~ \frac{ cm ^{3} }{ dm ^{3} }\] Mathmatical axiom that allow us to multiply with 1: \[1*x=x ~ ~ ~ \forall x\]

OpenStudy (anonymous):

Are you checking your significant figures?

OpenStudy (anonymous):

I've tried that as well and it is still wrong

OpenStudy (frostbite):

Well still weird. Think it is a bug of some kind.

OpenStudy (anonymous):

Try 7000

OpenStudy (anonymous):

If you can that is*

OpenStudy (anonymous):

That was it! Thanks

OpenStudy (anonymous):

Awesome. Remember, because 7 was a value, your significant figures became 1. so 7039 would become 7000.

OpenStudy (frostbite):

Emmm 7000 have 4 signaficant figures, the zeros too are significant.

OpenStudy (frostbite):

If you want only 1 significant figure we should say \[7*10^{3}\]

OpenStudy (anonymous):

But I do not have a period after the 7000 So it has 1 sig-fig ;o